OpenStudy (anonymous):

The equation of a line is y = 2x + 3. What is the equation of the line that is parallel to the first line and passes through (2, –1)?

OpenStudy (tkhunny):

Please write out the Point-Slope form of a line.

OpenStudy (anonymous):

A - 4x – 2y = –6 B. y = 2x – 5 C. y = 3x + 4 D. 2x + y = –1 one of these are the answer

OpenStudy (tkhunny):

Oh, well, in that case, Please write out the Point-Slope form of a line.

OpenStudy (anonymous):

two paralell eq has same slope..so slope is=2 and from point (2,-1) use point slope form ..i.e. (y-y1)=m(x-x1) solve it and get result..

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ y = 2x + 3 }\) \(\normalsize\color{blue}{ y = 2x + b }\) \(\normalsize\color{blue}{ (-1) = 2(2) + b }\) \(\normalsize\color{blue}{ -1 = 4 + b }\) \(\normalsize\color{blue}{ -5 = b }\) \(\normalsize\color{blue}{ y = 2x - 5 }\)